JoelTheBoss (joel_the_boss):

Write the equation of the line which passes through (1, 2) and with slope of − 1 3 in standard form. Then Solve.

JoelTheBoss (joel_the_boss):

@Walkerman15 if you dont get it its ok

JoelTheBoss (joel_the_boss):

do you get it?

OpenStudy (yanasidlinskiy):

What's the question, really? I'm not understanding it.

JoelTheBoss (joel_the_boss):

you have to get into y=mx+b form then you have to solve it :3

JoelTheBoss (joel_the_boss):

x+3y = 7 I have the form I just need to solve it now :( @Walkerman15 @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

The first thing we do..Is...-x from both sides..Re-arrange the equation!! -3y=-x+7.

JoelTheBoss (joel_the_boss):

I don't get any of this stuff can you explain :(

OpenStudy (yanasidlinskiy):

You're trying to get y by itself. Correct?

JoelTheBoss (joel_the_boss):

YES

ganeshie8 (ganeshie8):

\[\large x+3y = 7 \] subtract -x both sides, you get \[3y =-x+ 7 \]

ganeshie8 (ganeshie8):

next divide 3 both sides

JoelTheBoss (joel_the_boss):

ok

JoelTheBoss (joel_the_boss):

x+y=2.33

ganeshie8 (ganeshie8):

\[\large \dfrac{3y}{3} =\dfrac{-x}{3}+ \dfrac{7}{3}\]

ganeshie8 (ganeshie8):

\[\large y =\dfrac{-1}{3}x+ \dfrac{7}{3}\]

ganeshie8 (ganeshie8):

is that the form you wanted ?

JoelTheBoss (joel_the_boss):

I think so it looks like other things ive been working on.

ganeshie8 (ganeshie8):

good, if u compare both the equations you will see : slope = m = -1/3 y intercept = b = 7/3

JoelTheBoss (joel_the_boss):

y=-1/3x+7/3?